A call center employs 1,000 agents. Every month 50 employees leave the company and 50 new employees are hired.

a. How long on average does an agent work for this call center?
  Average working time months

Suppose the cost of hiring and training a new agent is $1,000. The manager of this call center believes that increasing agents’ salaries would keep them working longer at the company. The manager wants to increase the average time that an agent works for the call center to 24 months, or two years.

b-1.

Use the calculated average working time at this call center to determine the current annual cost for hiring and training.

  Current annual cost $
b-2.

Determine the new annual cost for hiring and training.

  New annual cost $
b-3.

How much can the company save on hiring and training costs over a year?

  Annual saving $   
Answer #1

a.

Throughput Rate = 50 people/month

WIP = 1000 people

Average working time = Throughput Time= WIP/TR = 1000/50 = 20 months = 1.67 years

b.

1. Current annual cost for hiring and training:

Throughput Rate = 50 people/month = 600 people/year

Annual hiring and training cost is 600*1000=$600,000

2. New annual cost for hiring and training:

Average working time = Throughput Time = 24 months = 2 years

Throughput Rate = WIP/TT = 1000 people/ 2 years = 500 people/ year

Annual hiring and training cost is 500*1000=$500,000

3. Annual saving on hiring and training cost is $100,000 ($600,000 - $500,000)